I had this question to answer and I'm not sure if the calcualtions are correct. Can someone give me some direction. Thanks
A ball drops from a certain height rises after bouncing to 40% of its original height. After bouncing if the ball rises to 2 ½ ft. What is the height from which it was dropped?
a = b x r x100
a = 2 ½ x 40% (Change 2 ½ to decimal and 40% to a decimal)
a = 2.5 x .40 x 100
a = 2/5 x .40 = 1 ft.

Answer by Alan3354(69443)

A ball drops from a certain height rises after bouncing to 40% of its original height. After bouncing if the ball rises to 2 ½ ft. What is the height from which it was dropped?
a = b x r x100
a = 2 ½ x 40% (Change 2 ½ to decimal and 40% to a decimal)
a = 2.5 x .40 x 100
a = 2/5 x .40 = 1 ft.
--------------------------
It would bounce to 1 foot on the next bounce. That's not what was asked for, tho.
-------------
The question is, where did it start when it bounced to 2.5 feet.
2.5 feet is 40% of the height it was dropped from.
0.4*h = 2.5
h = 2.5/0.4
h = 6.25 feet